### v1

# Perception-Distortion Balanced Super-Resolution: A Multi-Objective Optimization Perspective

Abstract
High perceptual quality and low distortion degree are two important goals in image restoration tasks such as super-resolution (SR). Most of the existing SR methods aim to achieve these goals by minimizing the corresponding yet conflicting losses, such as the $\ell_{1}$ loss and the adversarial loss. Unfortunately, the commonly used gradient-based optimizers, such as Adam, are hard to balance these objectives due to the opposite gradient decent directions of the contradictory losses. In this paper, we formulate the perception-distortion trade-off in SR as a multi-objective optimization problem and develop a new optimizer by integrating the gradient-free evolutionary algorithm (EA) with gradient-based Adam, where EA and Adam focus on the divergence and convergence of the optimization directions respectively. As a result, a population of optimal models with different perception-distortion preferences is obtained. We then design a fusion network to merge these models into a single stronger one for an effective perception-distortion trade-off. Experiments demonstrate that with the same backbone network, the perception-distortion balanced SR model trained by our method can achieve better perceptual quality than its competitors while attaining better reconstruction fidelity. Codes and models can be found at https://github.com/csslc/EA-Adam.
Index Terms:
Super-resolution, GAN, evolutionary algorithm, optimization.
I Introduction
Image super-resolution (SR) [1] is a valuable technique to reconstruct a high-resolution (HR) image with good quality from a low-resolution (LR) input image. Benefiting from the deep learning techniques, it has become popular to train a deep neural network (DNN) for SR by using a large amount of LR-HR image pairs [2, 3, 4, 5, 6, 7, 8]. One commonly used loss function to optimize the SR models is the distortion (or fidelity) measure, e.g., the norm of the prediction errors [2, 9, 10, 11]. However, SR is a typical ill-posed problem, and there are many possible HR estimations for the given LR input. It is well-known that the $\ell_{2}$ or $\ell_{1}$ loss tends to generate over-smoothed HR estimates [12, 13] although they can result in good distortion measures such as peak-to-noise ratio (PSNR).
To better preserve the local structures and details of reconstructed images, the SSIM loss [14] and perceptual loss [15] have also been used to train SR models. The former adopts the widely used image quality assessment index, i.e., SSIM [14, 16], as the optimization objective, while the latter minimizes the $\ell_{1}$ or $\ell_{2}$ norm of the prediction difference in a feature space [7, 8]. In general, the SSIM loss and perceptual loss can improve the perceptual quality of SR outputs, but the improvement is limited because they can hardly generate details that are lost during degradation[17, 18].
Figure 1: Performance comparison of perception-distortion balanced SR models trained by the Adam optimizer and our hybrid EA-Adam optimizer on the BSD100 dataset. For the perceptual quality index LPIPS, the smaller the better. For the distortion degree index PSNR, the higher the better. The models optimized by Adam (orange triangle points) with different weighted combinations of $\ell_{1}$, perceptual and adversarial losses are either stagnated into an unsatisfactory solution, or achieve unbalanced performance. The models optimized by our EA-Adam optimizer (blue circle points) can obtain better convergence and divergence. Furthermore, the fused model (red star) from EA-Adam models achieves the best perceptual quality index while maintaining a high fidelity index.
To obtain perceptually more realistic HR images, the perception-distortion trade-off methods [19, 7, 8, 20, 21, 22] have been developed in SR. Usually, they train a generative adversarial network (GAN) [23] by using the adversarial loss. Instead of minimizing the pixel-wise $\ell_{2}$ or $\ell_{1}$ loss or local structure-wise SSIM or perceptual loss, the adversarial loss is optimized to minimize the distribution divergence between the predicted SR images and HR images. In practice, it is unstable and even impossible to train the SR model by using only the adversarial loss. Therefore, most of the GAN-based SR methods weigh the $\ell_{1}$, perceptual and adversarial losses [7, 8] into one single loss for optimization.
While GAN-based SR methods can generate a certain amount of realistic details, they will also introduce undesired artifacts [24, 17]. Many following works [25, 12, 26, 27, 28] aim to reduce the artifacts caused by the adversarial loss while keeping the realistic details. Dario et al[25] applied $\ell_{1}$ and adversarial losses in both spatial and Fourier domains to improve the restoration of high-frequency details. Liang et al[12] designed an LDL loss to suppress GAN-generated artifacts by considering their statistical difference from visually friendly details. To avoid artifacts in low-frequency areas, a region-aware adversarial loss [26] was designed to help the discriminator pay more attention to high-frequency areas. Zhang et al[27] used a similar idea to handle the low-frequency and high-frequency areas differently by using a constrained loss. Xie et al[28] proposed to detect the artifact regions and developed a finetuning procedure to improve GAN-based SR models.
Unfortunately, the above-mentioned methods have two major limitations. First, empirically selecting weights to combine the losses into one is unreasonable and labor-consuming [27]. Second and more importantly, distortion- and perception-oriented losses often conflict with each other, and the commonly used gradient-based DNN optimizers, such as Adam [29], are difficult to balance the conflicting gradient-decent directions. We illustrate this issue in Fig. 1, where we train a population of SR networks by weighing the $\ell_{1}$, perceptual and adversarial losses with different weights. The Adam optimizer is used to train the network, and the PSNR (the higher the better) and LPIPS (the smaller the better) [30] indices are used to measure the distortion degree and perceptual quality, respectively. An ideal SR model is expected to achieve both high PSNR and low LPIPS, i.e., approaching the origin point in Fig. 1. However, we can see that the learned models (orange triangle points) by Adam with different weighted $\ell_{1}$, perceptual and adversarial losses are rather stagnated into an unsatisfactory solution, or achieve unbalanced performance.
In this paper, we propose a new approach for perception-distortion balanced SR from the network optimization perspective. We formulate the problem as a multi-objective optimization task [17], one objective aiming for high perceptual quality and another for low distortion degree. Considering that these two objectives are hard to achieve simultaneously by using gradient-based optimizers because of their conflicting gradient descent directions, we propose to integrate the gradient-free evolutionary algorithm (EA) [31, 32, 33] with the gradient-based Adam for network optimization. Adam mainly ensures the model convergence, i.e., the losses will decrease along the gradient, while EA mainly ensures the model divergence, i.e., the potential solutions could interact with and benefit from each other by the crossover operation. Meanwhile, the mutation operation of EA helps the model to jump out of the local optimum. As a result, the obtained Pareto front (PF) consists of a series of optimal models with different perception-distortion preferences, as shown by the blue circular points in Fig. 1. Compared with the Adam optimizer, the models optimized by our proposed EA-Adam method are closer to the original point and are more evenly distributed than those optimized by Adam. Motivated by the mixture of experts theory [34, 35] and the network interpolation methods [36, 37], we further propose a simple yet effective fusion network to merge the obtained models into a single stronger one, which inherits the advantages of the model population. As indicated by the red star in Fig. 1, the fused model achieves the best perceptual quality index while maintaining a high fidelity index. Extensive experiments on benchmark datasets demonstrate the effectiveness of the proposed hybrid EA-Adam optimizer and weight aggregation network. Compared with the existing state-of-the-art methods, our model achieves a better perception-distortion trade-off.
The rest of this paper is organized as follows. Section II summarizes the related works on distortion- and perception-oriented SR methods, and evolutionary algorithms on DNN learning. Section III presents the multi-objective formulation of SR, the hybrid EA-Adam optimizer, and the network fusion approach in detail. Section IV reports the experimental results and discussions. Finally, Section V concludes this paper.
II Related Work
II-A Distortion-oriented SR Methods
Most of the SR methods in this category [2, 3, 10, 5] employ the pixel-wise loss functions, such as the $\ell_{2}$ or $\ell_{1}$ norm of the prediction error, while focusing on the design of network architectures to improve the SR performance. The seminal work of SRCNN [2] introduced a convolutional neural network (CNN) with three layers to perform the SR task. Afterward, many SR models have been developed. Kim et al[6] trained a deeper and wider network to improve the representational ability of CNN. Deep residual connections were introduced to improve deep SR network learning performance [38, 39, 9]. Tong et al[11] proposed to use dense skip connections in a very deep network. The recursive network et al[6] was designed to improve SR performance without introducing new parameters for additional convolutions. Very recently, the transformer networks [40, 41, 42, 43] have been successfully employed for SR, resulting in better distortion-based measures such as PSNR. Because $\ell_{2}$ or $\ell_{1}$ loss tends to generate blurry SR outputs, the SSIM loss [14] and perceptual loss [15] have been proposed to regularize the local structures of images. These two losses can improve the visual quality of SR outputs without harming too much the distortion measures. However, they cannot generate image details to make the SR image perceptually more realistic.
II-B Perception-oriented SR Methods
To make the reconstructed SR images look more perceptually realistic, SRGAN [7] combined the adversarial loss with distortion-oriented losses (e.g., $\ell_{1}$ loss) in training. Following SRGAN, ESRGAN [8] used the VGG features before activation and employed the RRDB backbone [8] as the generator model, improving much the perceptual quality. While SRGAN and ESRGAN can generate new image details, they introduce many unnatural visual artifacts because of the unstable adversarial training. Therefore, many of the following works aim to reduce the GAN-generated artifacts from the perspectives of frequency division [26, 25, 27, 44], or different statistical properties [12, 28]. In addition, some works [21] aimed to optimize the no-reference image quality metrics, such as NIQE [45] and Ma [46], which are non-differentiable. For example, RankSRGAN [21] designed a differentiable rank-content loss to rank the quality of predictions with NIQE or Ma index.
II-C Evolutionary Algorithms in DNN Learning
EA [32, 33] is a classical search-based optimization technique, where individuals interactively evolve from a population. The EA operators mainly consist of crossover [47], mutation, and selection. Due to the global search and gradient-free properties, EA has proven its effectiveness in many applications [48, 49, 50]. Recently, it has also been applied to DNN learning [51, 52, 53]. Guo et al[51] applied evolutionary architecture search to construct a simplified supernet with all architectures optimized simultaneously. Yu et al[52] searched over a big single-stage model that contains small children models. EAGAN [53] applied an efficient two-stage EA-based network search framework to learn GANs. Another application is network compression. For example, Zhou et al[54] established filter pruning as a multi-objective optimization problem and proposed a knee-guided EA algorithm to trade-off between the scale of parameters and performance. In addition, EA algorithm can be used as the network optimizer. ESGD [55] alternately applied SGD and EA in one framework to improve the average performance of DNN. Zhang et al[56] proposed a hierarchical cluster-based suppression algorithm to remove similar weights and improve population diversity. GEMONN [57] determined the search direction in the EA optimizer by the gradient of weights to improve the efficiency of EA for training DNN.
To optimize a network, EA is usually coupled with a gradient-based optimizer to solve large-scale optimization problems [58]. Most of the previous EA-based DNN learning works [55] optimize a single objective. In this work, we formulate the perceptually realistic SR task as a multi-objective optimization problem and propose a hybrid EA and Adam algorithm to solve it.
III The Proposed Method
TABLE I: {justify} The proposed hybrid EA-Adam optimization algorithm.
Input: The number of optimized models $N$, Adam epochs $T^{Adam}$, EA epochs $T^{EA}$, total epochs $T$, the pretrained model ${\theta}_{G_{0}}$ optimized by $\ell_{1}$ loss, and the probability $\delta$ of selecting parents from the population $\mathcal{B}$.
Step 1. Initialization: Initialize a population of generators $\mathcal{G}=\{\theta_{G_{1}},\cdots,\theta_{G_{N}}\}$, where ${\theta}_{G_{k}}={\theta}_{G_{0}},k\in[1,\cdots,N]$. Initialize the discriminator population ${\mathcal{D}}=\{\theta_{D_{1}},\cdots,\theta_{D_{N}}\}$. Generate the neighboring population $\mathcal{B}$. Define the Adam optimizers for each model.
Step 2. Optimization iteration:
For $t\in 1:T^{EA}+T^{Adam}:T$ do
##\##Adam Steps
For $i\in\{1,\cdots,T^{Adam}\}$ do
For $s\in\{2,\cdots,N\}$ do
1. Update the weights of $\theta_{G_{s}}$ by Adam.
2. Update the weights of $\theta_{D_{s}}$ by Adam.
EndFor
EndFor
##\##EA Steps
For $j\in\{1,\cdots,T^{EA}\}$ do
For $k\in\{2,\cdots,N\}$ do
1. If $r<\delta$, $\mathcal{P}=\mathcal{B}(k)$; otherwise, $\mathcal{P}=\mathcal{G}$.
2. Select parents ${\theta}_{1}$ and ${\theta}_{2}$ from $\mathcal{P}$ randomly.
3. Generate offspring ${\theta}_{I}$ by the crossover and mutation in Eq. (3) and Eq. (4).
4. Select the new individual $\theta_{G_{k}}$ as in Eq. (5).
EndFor
EndFor
Re-initialization: Re-define the Adam optimizer for each model in population.
EndFor
Figure 2: (a) The framework of traditional GAN-based SR methods. (b) The Adam steps optimize a set of generator and discriminator networks. (c) The EA steps optimize the generator networks by fixing the discriminator networks.
This section presents the proposed hybrid EA-Adam optimizer and a simple yet effective fusion network for perception-distortion balanced SR. Firstly, we formulate the SR task as a multi-objective optimization problem, with two objectives focusing on perceptual quality and fidelity quality, respectively. Secondly, a hybrid EA-Adam optimizer is designed to solve the problem, where EA addresses the model divergence and Adam addresses the model convergence. A set of SR models with different perception-distortion preferences are trained to form a PF. Lastly, a simple yet effective fusion network is proposed to merge the learned models into a strong perception-distortion-balanced SR model.
III-A Multi-Objective Formulation
Most of the GAN-based SR methods [7, 8] train the generator model by weighting the pixel-wise $L_{pix}$ loss, perceptual loss $L_{percep}$ and adversarial loss $L_{adv}$ as follows:
${L_{GAN}}={\alpha_{1}}{L_{pix}}+{\alpha_{2}}{L_{percep}}+{\alpha_{3}}{L_{adv}},$ (1)
where the $L_{pix}$ (i.e., $\ell_{1}$ loss) optimizes the reconstruction fidelity, and $L_{percep}$ and $L_{adv}$ improve the perceptual quality. The weights $\alpha_{1}$, $\alpha_{2}$ and $\alpha_{3}$ are often empirically set to 0.01, 1 and 0.005, respectively.
The Adam optimizer [29] is dominantly used to optimize the above loss function. Unfortunately, the gradient-based Adam is difficult to minimize the perception- and distortion-oriented losses simultaneously due to their opposite gradient descent directions, while an ideal weight setting is hard to be found given an infinite searching space. Therefore, most traditional SR models turn to pursue either fidelity or perceptual quality yet sacrifice the other. As illustrated in Fig. 1, when the weights are biased towards distortion- or perception-oriented losses, the SR results will be over-smoothed or include visually unpleasant artifacts. When similar weights are set to balance perception and distortion, the optimization might be stagnant at the local minimum as no explicit gradient descent directions can be found. Note that we apply gradient normalization [59] to avoid the trivial solution.
In this paper, rather than linearly combining the distortion- and perception-oriented losses into a single objective function, we formulate the perceptually realistic SR task as a multi-objective optimization problem as follows:
$\left\{\begin{array}[]{l}\min{f_{1}}={L_{pix}},\\ \min{f_{2}}={L_{percep}}+{\alpha}{L_{adv}},\end{array}\right.$ (2)
where objectives $f_{1}$ and $f_{2}$ respectively focus on the reconstruction fidelity and the perceptual quality of the SR results. In the following, we propose to couple EA [33] with Adam [29] to address the multi-objective optimization problem in Eq. (2). EA and Adam help with the optimization process for the divergence and convergence, respectively.
III-B The Hybrid EA-Adam Optimizer
The algorithm of the proposed hybrid EA-Adam optimizer is summarized in Table I. EA and Adam alternately optimize the $N$ SR models with different perception-distortion preferences. We employ the general framework of GAN-based SR learning, which is shown in Fig. 2(a). The generator network learns the mapping between the LR and HR images, and the discriminator network discriminates between the true and fake images. The two networks compete with each other to ensure that the generated images are of high quality and more natural.
The Adam step is illustrated in Fig. 2(b). Initialized by a pre-trained model with $\ell_{1}$ loss, the other $N-1$ generator-discriminator pairs are updated by Adam separately, as shown in the Adam step of Table I. We fix the first model across the training process as it has the best reconstruction fidelity by training with pure pixel-wise $\ell_{1}$ loss. Then, the EA step starts, as illustrated in Fig. 2(c). Note that updating the discriminator networks in EA or not has little impact on the final SR performance. This may be because the EA step focuses on the model divergence, while the adversarial training facilitates the model convergence in the Adam step. Therefore, the discriminator models are only updated in Adam steps to improve training efficiency.
In the EA step, a population of generator models is optimized together. Following MOEA-D [33], the parents are first chosen from the neighboring population $\mathcal{B}$ or the whole population $\mathcal{G}$, which focuses on the local or global evolution. The offspring is generated by the selected parents via crossover and mutation operators. The crossover is used to interact with the individuals in a population with each other so that useful information can be effectively propagated internally. We employ the SBX crossover [47], which is defined as follows:
${\theta}_{I}=0.5\times\left[{(1+\beta){\theta}_{1}}+(1-\beta){{\theta}_{2}}\right].$ (3)
Here, $\{\theta\}_{I,1,2}$ indicate the parameters of the corresponding models, $\beta$ is a weighting variable that relies on a random number $r\in[0,1]$, where $\beta={{{(r\times 2)}^{1/(1+\eta)}}}$ if $r<0.5$ and $\beta={{{(1/(2-r\times 2))}^{1/(1+\eta)}}}$ otherwise, $\eta$ is a constant and is set to 20 as in SBX [47]. The mutation operator disturbs the optimization to jump out of the local optimum by adding a zero-mean random Gaussian noise ${\varepsilon^{(k)}}\sim{\mathcal{N}(0,0.01)}$ [55] to each of the $N$ models, i.e.,
${\theta_{I}}={\theta_{I}}+{\varepsilon^{(k)}},k\in\{1,\cdots,N\}.$ (4)
To select a good model from the parent models and the newly-generated offspring models, we measure the model performance by using an aggregation function derived from the Tchebycheff decomposition [60] as follows:
$\min{F^{te}}=\max\{\lambda({f_{1}}-z_{1}^{*}),(1-\lambda)({f_{2}}-z_{2}^{*})\},$ (5)
where $z_{1}^{*}$ and $z_{2}^{*}$ are the moving minimum values of $f_{1}$ and $f_{2}$ recorded until the current iteration, and weights $\lambda$ are uniformly distributed within $(0,1]$. Specifically, we have $\lambda={0,\frac{1}{{N-1}}},{\frac{2}{{N-1}},\cdots,1}$ for the $N$ models, respectively. Note that the momentum in the Adam optimizer should be initialized after the EA optimization step. Finally, a set of generator models with different perception-distortion preferences are obtained to form a PF, as shown in Fig. 1. We will see that the SR models optimized by the hybrid EA-Adam optimizer can achieve both better perceptual quality and reconstruction fidelity than the corresponding models optimized by Adam, as validated in Section IV-C.
Figure 3: Two stages of our network fusion process. Left: training of the attention-based weight regression network, which predicts the fusion weights of experts for each input image. Right: model fusion by averaging weight vectors of a batch of validation data.
III-C Network Fusion
As mentioned above, the $N$ SR models (with the same architectural topology) optimized by the proposed hybrid EA-Adam optimizer can work better than their counterparts optimized by Adam in balancing perceptual quality and distortion degree. However, it can be labor-consuming to manually select one from them to super-resolve the input LR image. On the other hand, if we apply all the $N$ models to the input image and then fuse the outputs into one image, it can be expensive and less effective in inference [36]. To solve this issue, we propose to automatically fuse the $N$ models into one stronger model to pursue further enhanced performance beneath the PF formed by the $N$ models. To achieve this goal, we first introduce a lightweight attention-based weight regression network to fuse the convolution layers of the $N$ SR models as in [35, 34, 61]. As shown at the left of Fig. 3, the input is the feature from one layer and the output is the weight of the corresponding layer in the $N$ models. The generated weights satisfy $\sum\nolimits_{k}{{w_{k}}=1},k\in\{1,\cdots,N\}$. There are 3 modules in the weight regression network. The first module consists of 3 convolution layers with Leaky ReLU activation. The kernel sizes of convolution layers are all $5\times 5$. The second module is a global average pooling layer, and the last is a mapping module, which consists of two linear layers followed by a Sigmoid activation. The first module extracts the features and the last two modules estimate the weights to fuse the convolution layers of the SR models.
We train the attention-based weight regression network with the most commonly-used $L_{GAN}$ loss as it encourages the output to be perceptually more realistic. Meanwhile, since the aggregated model linearly interpolates among the $N$ SR models that are fixed in this stage, the reconstruction fidelity can also be well preserved. We fuse the $N$ SR models in one shot and apply the fused model to all testing images, as shown at the right of Fig. 3. Here, we employ a set of $M$ validation images, i.e., $\{I_{i}\}_{i=1}^{M}$, to facilitate the fusion process. Based on the validation data, we calculate the adaptive fusion weights for each $I_{i}$, i.e., $[w_{1}^{i},w_{2}^{i},\cdots,w_{N}^{i}]$, using the trained attention-based weight regression network, and average them over the $M$ weight vectors to get a universal weight, i.e., $[\bar{w}_{1},\bar{w}_{2},\cdots,\bar{w}_{N}]$. Finally, a fused model is obtained, which inherits the advantages from the $N$ SR models in terms of both reconstruction fidelity and perceptual quality, while holding a stronger generalization capability.
IV EXPERIMENT
IV-A Experiment Setup
Datasets and evaluation metrics. We use the DIV2K dataset [62] to train the population of SR models. We use 700 images in DIV2K for training the weight regression network and the remaining 100 images as the validation data in the network fusion process. We evaluate the performance of different methods on 6 benchmarks, including Set5 [63], Set14 [64], BSD100 [65], Urban100 [66], Manga109 [67], and DIV2K100 [62]. We compute the PSNR and SSIM [14] indices on the Y channel in the YCbCr space to measure the distortion degree and compute the LPIPS [30] index in the RGB space to evaluate the perceptual quality.
Backbones and compared methods: As in many previous perceptually realistic SR works [21, 20, 12, 26], we adopt SRResNet [7] and RRDB [8] as the backbone networks. SRResNet is a lightweight network proposed in SRGAN [7], while RRDB [8] is widely used in later GAN-based SR methods [20, 68, 12] for its superior performance. We compare our two models with the methods using the SRResNet and RRDB backbones, respectively. Specifically, we compare our SRResNet-based model with SRGAN [7] and RankSRGAN [21], and compare our RRDB-based model with ESRGAN [8], USRGAN [20], SPSR [68], LDL [12] and CAL-GAN [69]. For the methods that do not use these two backbones, we compare them with our model of similar capacity. Specifically, we compare our SRResNet-based model with CFSNet [70] and G-MGBP [19], and compare our RRDB-based model with EPSR[71] and PD-ADMM [27]. In addition, we compare SwinIR [40] trained with the $L_{GAN}$ loss by Adam optimizer and our proposed hybrid EA-Adam optimization method to validate the effectiveness of our method on the Transformer backbone. The results of the compared methods are obtained by using the officially released models. Note that we do not compare our method with the distortion-oriented SR methods, because this work focuses on perception-distortion balance.
Training details: We validate the effectiveness of our method with scaling factor 4 following the previous works [7, 8]. The data augmentation and discriminator networks used in SRGAN and ESRGAN are adopted in our SRResNet- and RRDB-based models, respectively, for a fair comparison. The size of training input patches is set as 32×32. For both the hybrid EA-Adam optimizer and the attention-based weight regression network, the learning rate is set as $1e^{-5}$ for the SRResNet backbone, and $1e^{-4}$ for the RRDB and SwinIR backbones. To balance the performance and the training efficiency, we set the population size as $N=5$. The $T^{Adam}$ for the Adam optimizer is 10, the $T^{EA}$ for the EA optimizer is 1, and the total optimization epoch $T$ is 100. The probability $\delta$ of selecting parents from the neighboring population $\mathcal{B}$ is 0.7.
TABLE II: {justify} Quantitative comparison between the state-of-the-art methods and the proposed hybrid EA-Adam optimization method. Two groups of methods trained with the same backbone are compared, where SRResNet [7] and RRDB [8] backbones are used, respectively. The best results of each group are highlighted in bold.
Method
SRGAN [7]
RankSR- GAN [21]
SRResNet + Ours
ESRGAN [8]
USRGAN [20]
SPSR [68]
LDL [12]
CAL- GAN [69]
RRDB + Ours
Set5
PSNR $\uparrow$
29.93
29.77
30.25
30.45
30.85
30.40
30.99
31.04
31.31
[…]
0.8535
0.8394
0.8584
0.8582
0.8627
0.8443
0.8679
0.8555
0.8792
[…]
0.0740
0.0763
0.0694
0.0741
0.0777
0.0686
0.0643
0.0670
0.0683
Set14
[…]
26.53
26.47
26.98
26.28
27.07
26.64
27.20
27.31
27.42
[…]
0.7173
0.7031
0.7342
0.7085
0.7324
0.7138
0.7429
0.7367
0.7537
[…]
0.1430
0.1398
0.1330
0.1315
0.1404
0.1345
0.1301
0.1308
0.1239
BSD100
[…]
25.58
25.51
25.94
25.33
26.03
25.51
26.12
26.29
26.51
[…]
0.6677
0.6510
0.6836
0.6643
0.6791
0.6599
0.6941
0.6829
0.7071
[…]
0.1770
0.1829
0.1766
0.1602
0.1747
0.1665
0.1614
0.1672
0.1609
Manga109
[…]
28.12
27.94
28.67
28.42
28.49
28.56
29.41
29.21
29.33
[…]
0.8656
0.8500
0.8735
0.8619
0.8658
0.8590
0.8767
0.8671
0.8874
[…]
0.0704
0.0755
0.0648
0.0644
0.0653
0.0664
0.0547
0.0690
0.0536
Urban100
[…]
24.41
24.53
24.83
24.36
24.83
24.80
25.50
25.33
25.57
[…]
0.7320
0.7284
0.7479
0.7363
0.7470
0.7473
0.7692
0.7639
0.7752
[…]
0.1439
0.1435
0.1354
0.1235
0.1321
0.1206
0.1096
0.1167
0.1150
DIV2K100
[…]
28.17
28.07
28.62
28.18
28.71
28.18
28.96
28.95
29.34
[…]
0.7765
0.7654
0.7900
0.7779
0.7902
0.7720
0.7970
0.7897
0.8123
[…]
0.1254
0.1318
0.1208
0.1151
0.1284
0.1126
0.1008
0.1072
0.1038
#FOLPs 10.38G 73.43G
#Params 1.52M 16.69M
TABLE III: {justify} Quantitative comparisons between the state-of-the-art perceptually realistic SR models and the models trained by our method. The ’RRDB+Ours-PD’ is based on the RRDB backbone with the weight setting in PD-ADMM [27]. The ‘SwinIR-$L_{GAN}$’ is the SwinIR model trained with the $L_{GAN}$ loss by the Adam optimizer. The best results of each group are highlighted in bold.
Method
G-MGBP  [19]
CFSNet [70]
SRResNet +Ours
EPSR  [71]
RRDB +Ours
PD-AD MM[27]
RRDB +Ours-PD
SwinIR -$L_{GAN}$ [40]
SwinIR +Ours
Set5
[…]
29.54
30.23
30.25
31.23
31.31 31.82
31.68
30.35
30.40
[…]
0.8419
0.8503
0.8584
0.8710
0.8792
0.8810
0.8838
0.8543
0.8581
[…]
0.0855
0.0781
0.0694
0.0771
0.0683
0.0758
0.0746
0.0694
0.0678
Set14
[…]
26.58
26.59
26.98 27.76
27.42
27.99
27.95
26.61
26.78
[…]
0.7151
0.7170
0.7342
0.7526
0.7537
0.7589
0.7668
0.7253
0.7302
[…]
0.1502
0.1619
0.1330
0.1518
0.1239
0.1401
0.1336
0.1268
0.1268
BSD100
[…]
25.68
25.60
25.94 26.85
26.51
26.86
26.93
25.58
25.75
[…]
0.6673
0.6677
0.6836 0.7076
0.7071
0.7137
0.7206
0.6764
0.6832
[…]
0.1882
0.1822
0.1766
0.2016
0.1609
0.1854
0.1774
0.1568
0.1560
Manga109
[…]
28.17
28.54
28.67 29.39
29.33
30.17
29.92
28.60
28.64
[…]
0.8609
0.8622
0.8735
0.8779
0.8874
0.8864
0.8952
0.8650
0.8694
[…]
0.0786
0.0689
0.0648
0.0736
0.0536
0.0616
0.0528
0.0631
0.0621
Urban100
[…]
24.40
24.72
24.83 25.72
25.57
26.29
25.89
25.06
25.14
[…]
0.7415
0.7445
0.7479
0.7703
0.7752 0.7882
0.7824
0.7564
0.7594
[…]
0.1477
0.1362
0.1354
0.1366
0.1150
0.1234
0.1220
0.1215
0.1193
DIV2K100
[…]
28.27
28.49
28.62 29.61
29.34
29.71
29.91
28.63
28.77
[…]
0.7766
0.7826
0.7900
0.8088
0.8123
0.8117
0.8226
0.7897
0.7945
[…]
0.1488
0.1233
0.1208
0.1329
0.1038
0.1223
0.1136
0.1044
0.1027
#FOLPs
22.42G
19.22G
10.38G
205.84G
73.43G
109.78G
73.43G
4.42G
4.42G
#Params
0.28M
5.00M
1.52M
43.09M
16.69M
18.67M
16.69M
0.93M
0.93M
Figure 4: Visual comparisons among the proposed EA-Adam optimization method and other state-of-the-art methods under the SRResNet backbone.
Figure 5: Visual comparisons among the proposed EA-Adam optimization method and other state-of-the-art methods under the RRDB backbone.
Figure 6: Visual comparisons among the proposed EA-Adam optimization method and PD-ADMM.
Figure 7: Visual comparisons among the SwinIR optimized with the $L_{GAN}$ loss by Adam optimizer and the proposed EA-Adam optimization method.
IV-B Comparison with State-of-the-Arts
In this section, we compare the final fused model by our proposed hybrid EA-Adam optimizer and network fusion method against state-of-the-art methods. Table II provides the quantitative comparisons among the methods with the same backbone. Table III provides the quantitative comparisons among the methods with similar model capacity.
From Table II, we can see that for the lightweight network models with the SRResNet backbone, SRGAN gets better PSNR and SSIM scores than RankSRGAN on almost all the benchmarks, while they have higher or lower LPIPS scores on different benchmarks. Our proposed method improves both the fidelity- and perceptual-oriented metrics by a large margin under all benchmarks without increasing the SR model complexity. This validates the advantages of our hybrid EA-Adam optimization strategy, which could push the PF toward the optimal point of solution space. For models trained with the RRDB backbone, we can see that our proposed method achieves better results than the competing methods ESRGAN, USRGAN, SPSR, LDL, and CAL-GAN in most cases. The LDL also demonstrates competing results in several cases, especially for the perceptual quality metric on the Urban100 dataset. This advantage comes from its iterative and local discriminative strategy that can effectively inhibit visual artifacts and encourage the generation of high-frequency details.
From Table III, we can see that the proposed method achieves consistent improvement on most benchmarks in terms of both perceptual quality (LPIPS) and reconstruction accuracy (PSNR and SSIM), although the compared methods G-MGBP, CFSNet, and EPSR have more model parameters and FLOPs (e.g., 22.42G FLOPs and 19.22G FLOPs for ‘G-MGBP’ and ‘CFSNet’, while 10.38G FLOPs for ‘SRResNet+Ours’). In addition, we test the released model of PD-ADMM, which performs favorably in terms of distortion control yet sacrifices the perceptual quality. Since PD-ADMM assigns equal weights, i.e., [0.5, 0.5], to the two objectives, we also apply this setting in Eq. (2). One can see that our method outperforms PD-ADMM in almost all cases but with fewer FLOPs and parameters. Furthermore, our method with the SwinIR backbone also demonstrates significant improvement over its counterparts, which verifies that our method is applicable to both CNN-based and transformer-based models.
We provide 4× SR visual comparisons between models (with SRResNet [7] and RRDB [8] backbones) trained by our method and other state-of-the-arts models in Fig. 4 and Fig. 5. We do not provide visual comparisons with the early published ESRGAN method due to limited space. We give the visual comparisons between PD-ADMM [27] and ‘RRDB+Ours-PD’, i.e., the RRDB model trained by our EA-Adam optimization method with the weight setting in PD-ADMM, in Fig. 6. The visual comparisons of models with the SwinIR [40] backbone are shown in Fig. 7. One can see that the proposed method can restore more correct visual patterns (e.g., the lines of zebra crossing in the third group of Fig. 5) and simultaneously inhibit the artifacts (e.g., the windows of building in the third group of Fig. 4). At the same time, our model generates many high-frequency details which facilitate perceptual quality (e.g., the lines of wood floor in the third group of Fig. 6).
IV-C Ablation Studies
To validate the effectiveness of the proposed hybrid EA-Adam optimization and network fusion methods, we conduct experiments to compare our hybrid EA-Adam algorithm with the Adam algorithm, and compare our network fusion method with the learnable weight method. Furthermore, different loss functions are applied to optimize the weight regression network to show the influence of loss selection. We also study the weights of obtained models with different perception-distortion preferences. All ablation results are evaluated on the Urban100 dataset.
TABLE IV: {justify} Quantitative comparison between the models optimized by Adam, AdamW, our hybrid EA-Adam, and hybrid EA-AdamW optimizers. The SRResNet backbone is adopted and the models are evaluated on the Urban100 dataset. ‘Fused Model’ is obtained by applying our network fusion method to the four trained models. The best results are highlighted.
Model
Model1
Model2
Model3
Model4
Fused model
Weights
$[0.25,0.75]$
$[0.5,0.5]$
$[0.75,0.25]$
$[1,0]$
PSNR
Adam
24.36
23.82
23.96
24.22
24.41
EA-Adam
25.78 25.75 24.95 24.73 24.83
AdamW
25.48
24.67
23.41
24.18
24.37
EA-AdamW
25.95 25.81 25.21 24.82 24.91
LPIPS
Adam
0.3061
0.2899
0.2665
0.1446
0.1437
EA-Adam
0.2120 0.2055 0.1536 0.1384 0.1354
AdamW
0.2570
0.2925
0.2306
0.1477
0.1444
EA-AdamW
0.2265 0.2201 0.1519 0.1404 0.1368
Figure 8: Visual comparisons among SR models trained by the original Adam optimizer and our hybrid EA-Adam optimizer with the SRResNet [7] backbone. Each model trained by EA-Adam outperforms the one trained by Adam in both fidelity and perceptual quality. The fused model by the proposed EA-Adam method achieves further benefits in fidelity and perception.
Hybrid EA-Adam vs. Adam. We compare the performance of models optimized by the original Adam and our hybrid EA-Adam. In addition, we replace Adam with AdamW [72], a recently developed variant of Adam, and report the results of AdamW and hybrid EA-AdamW for a more comprehensive evaluation of our proposed method. As traditional optimizers (Adam and AdamW) cannot directly optimize a multi-objective problem, we uniformly set 4 combinations of loss weights, i.e., $[0.25,0.75]$, $[0.5,0.5]$, $[0.75,0.25]$ and $[1,0]$, to weigh the two objectives in Eq. (2) into a single-objective optimization problem. For example, if the weight is $[0.25,0.75]$, the optimization objective of Adam is $0.25\times f_{1}+0.75\times f_{2}$, which is the same as that of the Adam steps in our proposed hybrid optimizer. The models optimized by different weights are named ‘Model1’ to ‘Model4’, respectively. In addition, we apply the same network fusion method (as in Section 3.3) to the 4 models optimized by Adam.
Table IV gives the quantitative comparison. Without loss of generality, we conduct experiments with the SRResNet backbone and evaluate on the Urban100 dataset. PSNR and LPIPS metrics are used to measure the fidelity and perceptual quality, respectively. One can see that our hybrid EA-Adam/AdamW optimizers outperform Adam/AdamW by a large margin in most cases. Both the fidelity and perceptual quality metrics are improved, validating that the proposed hybrid EA-Adam strategy can push the PF closer to the optimal point of the metric space. For most models, EA-Adam improves PSNR by more than 0.4dB over Adam, and improves LPIPS by more than $10\%$. This might be caused by the gradient vanishing problem in the gradient-based optimization process so that the models can be trapped into a local minimum. In contrast, with our hybrid EA-Adam optimizer, effective interaction between different models can be introduced by the crossover operator in EA, so that the models can be stably optimized with better convergence.
On the other hand, each of the four models may perform well in one aspect, yet sacrifices the other aspect as a trade-off. The fused model by our network fusion method achieves a better balance between fidelity and perceptual quality. The LPIPS of the fused model is even better than Model 4 trained to maximize perceptual quality. It is worth noting that the bad convergence of Models 1 to 3 makes the traditional Adam/Adamw optimizer gain less improvement from the network fusion process. In contrast, the proposed hybrid optimization strategy facilitates the effectiveness of network fusion via the interaction of individuals in crossover operation. In addition, we find that AdamW does not show much advantages over Adam in the SR task. This is because AdamW is mainly proposed to handle the over-fitting problem in high-level vision tasks by introducing weight decay. However, this may not fit the SR task as the optimization might be under-fitting to the large space of image details.
Fig. 8 shows the visual comparisons among SR models trained by Adam and our EA-Adam optimizers, where consistent observation with Table IV can be drawn. Specifically, models optimized by EA-Adam outperform their counterparts optimized by Adam, where the structures and patterns are more loyal to the HR and richer details are generated. It can also be observed that while each of the four models is preferred in either fidelity or perceptual quality, the fused model by our method shows comprehensive enhancement in both aspects.
TABLE V: {justify} Quantitative comparison between the ‘Learnable Weight’, ‘Weight Regression’, and ‘Averaging Fusion’ on the Urban100 dataset. The RRDB backbone is adopted. The best and second results are highlighted in red and blue.
PSNR
SSIM
LPIPS
Learnable Weight
25.60
0.7748
0.1195
Weight Regression
25.71 0.7756 0.1124
Averaging Fusion
25.57
0.7752 0.1150
Network Fusion vs. Learnable Weight. We propose a two-stage framework for fusing $N$ networks. The first stage is the attention-based weight regression, which generates a set of weights for each input LR image. Therefore, the corresponding weight vectors of validation data can be obtained. The second stage is model fusion, which obtains a single stronger model by averaging all the weight vectors of validation data. To show the effectiveness of the proposed fusion method, we introduce a fusion method by setting the fusion weights of $N$ models as learnable parameters directly, which is named ‘Learnable Weight’. We compare our two stages, named ‘Weight Regression’ and ‘Averaging Fusion’, with ‘Learnable Weight’ on the Urban100 dataset in Table V. We can see that the ‘Weight Regression’ achieves the best results in all metrics since it can dynamically infer the optimal weights for each input based on image content. The ‘Averaging Fusion’ obtains better performance than ‘Learnable Weight’ because the universal network weight is inherited from ‘Averaging Fusion’ using validation data, which holds a stronger generalization capability.
TABLE VI: {justify} Quantitative comparison on Urban100 benckmark among different loss functions for the network fusion process. The RRDB backbone is adopted.
PSNR
SSIM
LPIPS
RRDB+Ours-$\ell_{1}$
26.55
0.8029
0.1955
RRDB +Ours-PD [65]
25.89
0.7824
0.1220
[…]
25.57
0.7752
0.1150
Loss Selection for Network Fusion. The fusion network aims to leverage the strength of $N$ models for perceptually more plausible SR results. Table VI gives the numerical comparisons on the Urban100 benchmark with $\ell_{1}$, $PD$ [27] and $L_{adv}$ loss functions. We choose the adversarial loss in this paper because it achieves the best LPIPS score, which implies perceptually more realistic outputs.
TABLE VII: {justify} Weights of two layers (L1 and L2) in the RRDB models obtained with different perception-distortion preferences given two inputs (I1 and I2), and the final fusion model (Ours).
L1
L2
I1
[0.0463, 0.0167, 0.0678, 0.3333, 0.5358]
[0.2348, 0.2084, 0.1426, 0.1452, 0.2689]
I2
[0.0263, 0.0144, 0.1014, 0.3307, 0.5272]
[0.3371, 0.1061, 0.0022, 0.1985, 0.3561]
Ours
[0.0965, 0.0733, 0.1429, 0.2629, 0.4243]
[0.3203, 0.1212, 0.0224, 0.1932, 0.3428]
TABLE VIII: {justify} Quantitative comparisons on Urban100 benchmark between different ways of weight combination. The RRDB backbone is adopted. The best results are highlighted in bold.
PSNR
SSIM
LPIPS
21.86
0.6546
0.2269
24.52
0.7573
0.1755
25.57 0.7752 0.1150
Weights of Obtained Models. To further show the advantage of our proposed method, we study on the weights of obtained models. Given two inputs (I1 and I2), the weights in two layers (L1 and L2) of the $5$ obtained RRDB models are shown in Table VII. The weights of the final fusion model ‘Ours’ are also given. It can be seen that the weights are neither evenly distributed among models nor concentrated on one or two models. The weight distribution varies in layers, and each model contributes to the final fusion. This demonstrates the effectiveness of our EA-Adam optimizer, which can simultaneously obtain models with different focuses. We also perform the ablation study on weight by (a) directly averaging the parameters of the $5$ networks, (b) applying the weights of one layer to all layers of the $5$ networks, and (c) our averaging fusion strategy. The results on Urban100 are shown in Table VIII. Our averaging strategy gets the best result, showing the effectiveness of our network fusion method.
TABLE IX: {justify} Comparison of the training time of our hybrid EA-Adam optimization method and the Adam method.
Method Adam EA-Adam optimizer Network Fusion
EA-iter
Adam-iter
Run-time/epoch
0.19h
4.76h
0.80h
0.63h
Total-epoch
250
8
92
50
Total-run-time
48h
112h
32h
IV-D Training and Inference Time
Training time. We first compare the training time of our hybrid EA-Adam optimization method (including the EA-Adam optimizer and the network fusion stage) and the original Adam optimizer by training an SR model with RRDB-backbone for ×4 super-resolution. Table IX provides the quantitative comparison of the training time. Specifically, the Adam optimizer costs 250 epochs and 48h to converge. Though hybrid EA-Adam costs a longer time in each epoch, it converges in much fewer epochs thanks to the effective interaction in the EA step. The hybrid EA-Adam optimization method totally costs 144h to finish the model training, which is about three times the training time of Adam. Most of the training time of hybrid EA-Adam is spent on the EA part. It is noted that there have been various parallel computation methods [73, 74] to speed up the EA process based on CUDA technology, showing great potential to accelerate our method in training.
Inference time. While spending more time in training, the proposed method does not introduce any extra computational and memory burden during the inference stage. The trained SR model has the same inference time as other SR models with the same backbone.
V Conclusion
In this paper, we proposed a hybrid EA-Adam optimization method to optimize the perception-distortion balanced image super-resolution (SR) models. We formulated the perception-distortion trade-off as a multi-objective optimization problem. A population of SR models were optimized by alternatively performing the Adam and EA steps, where the Adam was dedicated to the model convergence and the EA focused on rescuing the model from local minimum due to its strong capability in multi-objective searching. An effective model fusion strategy was then designed to merge the trained models into a stronger one, which further improved the perception-distortion trade-off without increasing the inference cost. Extensive experiments demonstrated the effectiveness of the proposed method against the traditional Adam optimizer, and the state-of-the-art perception-distortion balanced performance of the trained SR models.